# IDESplat: Iterative Depth Probability Estimation for Generalizable 3D Gaussian Splatting

### v2

@@ Abstract
Generalizable 3D Gaussian Splatting aims to directly predict Gaussian parameters using a feed-forward network for scene reconstruction. Among these parameters, Gaussian means are particularly difficult to predict, so depth is usually estimated first and then unprojected to obtain the Gaussian sphere centers. Existing methods typically rely solely on a single warp to estimate depth probability, which hinders their ability to fully leverage cross-view geometric cues, resulting in unstable and coarse depth maps. To address this limitation, we propose IDESplat, which iteratively applies warp operations to boost depth probability estimation for accurate Gaussian mean prediction. First, to eliminate the inherent instability of a single warp, we introduce a Depth Probability Boosting Unit (DPBU) that integrates epipolar attention maps produced by cascading warp operations in a multiplicative manner. Next, we construct an iterative depth estimation process by stacking multiple DPBUs, progressively identifying potential depth candidates with high likelihood. As IDESplat iteratively boosts depth probability estimates and updates the depth candidates, the depth map is gradually refined, resulting in accurate Gaussian means. We conduct experiments on RealEstate10K, ACID and DL3DV. IDESplat achieves outstanding reconstruction quality and state-of-the-art performance with real-time efficiency. On RE10K, it outperforms DepthSplat by 0.33 dB in PSNR, using only 10.7% of the parameters and 70% of the memory. Additionally, our IDESplat improves PSNR by 2.95 dB over DepthSplat on the DTU dataset in cross-dataset experiments, demonstrating its strong generalization ability. The code is available at here.
@@ 1 Introduction
Efforts have been made to obtain accurate and refined depth estimation. Early methods [2, 34] directly use differentiable operations to predict the depth probability distribution from image features. Although these methods can predict the depth map for scene reconstruction in a generalizable way, their ability to exploit multi-view feature similarity is limited, which restricts their performance. Subsequent approaches [5, 21] introduce cost volumes that use warp operations to establish feature similarity across views, which provide valuable geometric cues for depth estimation and simplify the network learning process. However, these methods rely solely on a single warp to model feature similarity, which prevents them from fully exploiting the rich geometric information, leading to unstable and coarse depth maps. Therefore, how to incorporate multiple warps to gradually leverage rich cross-view geometric details, producing refined and reliable depth maps for accurate Gaussian mean prediction, remains a key challenge.
In this paper, we propose IDESplat, which iteratively performs warps to refine depth maps for accurate Gaussian means prediction. By integrating cascade warp results, we can progressively boost the feature similarity measure to identify high-likelihood surface points and suppress low-probability depth candidates. The iterative warp framework of our IDESplat is shown in Fig. 1. Firstly, to eliminate the instability of a single warp, we introduce a Depth Probability Boosting Unit (DPBU) to fuse multiple epipolar attention maps in a multiplicative manner for a reliable depth map. Next, we gradually update the depth search range while increasing the feature resolution, enabling warp and correlation calculations at a finer scale. Feature matching becomes easier and more precise as the depth search range is re-centered and image features are enhanced in this process. Finally, for other Gaussian parameters, we propose a Gaussian Focused Module that determines the most relevant Gaussian tokens to compute attention weights for feature interaction. The experimental results on RealEstate10K show that IDESplat outperforms DepthSplat by 0.33 dB in PSNR with only 10.7% of the parameters and 70% of the memory. Moreover, IDESplat shows outstanding cross-dataset generalization, achieving 28.79 dB on ACID when transferred from RE10K, outperforming methods trained directly on ACID. It also improves PSNR by 2.95 dB over DepthSplat on the DTU dataset in cross-dataset experiments. In summary, the main contributions of this paper are as follows:
@@
  • To eliminate the inherent instability of a single warp, we introduce a Depth Probability Boosting Unit that multiplicatively integrates multiple epipolar attention maps for reliable and refined depth map estimation.
@@ 2 Related Work
Generalizable 3D Gaussian Splatting. Single-scene 3D Gaussian Splatting methods [17, 9, 47, 28, 10] achieve more efficient rendering than neural fields and volume rendering methods [23, 31, 26, 40] due to their rasterization-friendly properties, but are still constrained by long optimization times and limited generalization capabilities. Generalizable 3D Gaussian Splatting methods [2, 35, 22, 46] address this issue through single-pass feed-forward prediction of all Gaussian parameters within a scene, enabling fast modeling of unseen scenes and view synthesis. Early studies [2, 34] used feed-forward neural networks to directly predict Gaussian parameters from image features, enabling generalizable scene reconstruction. Since the introduction of cost volume computation to generalizable 3DGS, many studies [5, 21] have focused on extracting geometric cues from cross-view images to enhance depth estimation accuracy, thereby simplifying the optimization of Gaussian parameters. Subsequently, recognizing the importance of accurate depth estimation, some methods [46, 22] have incorporated pre-trained monocular depth models, leading to significant performance improvements. However, Existing methods rely on a single warp for depth estimation, limiting their ability to exploit cross-view feature cues. In contrast, our IDESplat integrates feature similarity from cascaded warps, producing reliable and refined depth maps for Gaussian mean prediction.
Iteration-based Optimization Methods. Iteration-based optimization methods [1, 18] are widely used in various tasks [13, 55, 3] due to their ability to progressively enhance and refine the feature learning process. For instance, in optical flow estimation [36, 14, 48, 33, 15], iterative optimization is commonly used to build coarse-to-fine pyramidal features by stacking multiple feature units, resulting in more accurate and stable flow estimation. In the field of depth estimation, some works [19, 37, 41, 42, 6] have iteratively retrieved multi-view correlation features using structures like GRUs and LSTMs with shared parameters to update the disparity field, yielding better depth estimation results. Similarly, in monocular depth estimation [56, 30, 11], iterative methods are often used to gradually refine the depth search range, leading to more stable and accurate predictions. Additionally, in the 3D Gaussian Splatting (3DGS) domain, recent methods [38, 52, 45, 4, 12] have also attempted to design iterative optimization processes to reduce the difficulty of 3D Gaussian reconstruction tasks. Unlike existing iterative optimization methods, we propose a novel Depth Probability Estimation Unit that integrates the similarity results of multiple warps in a multiplicative manner. This approach produces more reliable and refined depth maps, while progressively enhancing feature resolution and refining the depth candidate range throughout the iterative process.
@@ 3 Method
Figure 2: The overall architecture of IDESplat. IDESplat is composed of three key parts: a feature extraction backbone, an iterative depth probability estimation process, and a Gaussian Focused Module (GFM). The iterative process consists of cascaded Depth Probability Boosting Units (DPBUs). Each unit combines multi-level warp results in a multiplicative manner to mitigate the inherent instability of a single warp. As IDESplat iteratively updates the depth candidates and boosts the probability estimates, the depth map becomes more precise, leading to accurate Gaussian means.
@@
Given a sequence of $V$ sparse-view images $\mathcal{I}=\{{\bm{I}}^{i}\}_{i=1}^{V}$, where each image ${\bm{I}}^{i}$ has dimensions $H\times W\times 3$, and the corresponding camera projection matrices are ${\bm{P}}^{i}\in\mathbb{R}^{3\times 4}$. The goal of the generalizable 3DGS task is to learn a network that maps images to 3D Gaussian parameters. This process is defined as:
$f_{\bm{\theta}}:\{({\bm{I}}^{i},{\bm{P}}^{i})\}_{i=1}^{V}\mapsto\{(\bm{\mu}_{j},\bm{\alpha}_{j},\bm{\Sigma}_{j},\bm{c}_{j})\}_{j=1}^{V\times H\times W},$ (1)
where $f_{\bm{\theta}}$ is a feed-forward network with learnable parameters $\bm{\theta}$}. The parameters $\bm{\mu}_{j}$, $\bm{\alpha}_{j}$, $\bm{\Sigma}_{j}$, and $\bm{c}_{j}$ are the Gaussian mean, opacity, covariance, and color, respectively. Since Gaussian means are difficult to predict directly, most existing methods estimate them by unprojecting a depth map into 3D. The depth estimation scheme is as follows: First, the input image sequence is processed using a multi-view feature extraction backbone, resulting in a downsampled feature $\bm{F}\in\mathbb{R}^{V\times\frac{H}{4}\times\frac{W}{4}\times C}$. Then, the cost volume is computed based on $\bm{F}$, the camera projection matrices ${\bm{P}}^{i}$, ${\bm{P}}^{j}$ for different views, and the depth candidates ${\bm{G}}=[d_{1},d_{2},\cdots,d_{D}]\in\mathbb{R}^{D}$. Specifically, the feature ${\bm{F}}^{j}$ from view $j$ is warped to view $i$ as follow:
${\bm{F}}^{j\to i}=\mathcal{W}({\bm{F}}^{j},{\bm{P}}^{i},{\bm{P}}^{j},\bm{G})\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times D\times C},$ (2)
@@
where $m\in\{1,\ldots,D\}$. Subsequently, $\bm{C}^{i}=[\bm{C}_{d_{1}}^{i},\ldots,\bm{C}_{d_{D}}^{i}]$ is refined with a U-Net and upsampled to the input resolution as $\tilde{\bm{C}}^{i}\in\mathbb{R}^{H\times W\times D}$. Finally, a $\mathrm{softmax}$ function is applied to obtain the probability of each candidate depth, and the final depth map is computed as their weighted average.
To achieve high-quality 3D reconstruction, an accurate depth map $\bm{D}$ is crucial since it determines the centers of 3D Gaussians. However, this commonly used method only relies on a single cross-view warp to compute feature similarity, underutilizing cross-view geometry and often leading to unstable, coarse depth estimates. In addition, this typical correlation computation method requires storing all the dense warping features, which leads to a memory issue, especially when the depth candidate size $D$ is large.
@@ 3.2 Iterative Multi-View Depth Estimation
Warp-Index Epipolar Attention. The warp operation is key to constructing the cost volume or epipolar attention map, as it establishes pixelwise correspondences between features in the source and target views. However, as shown in Eq. (2), this warping operation requires sampling target-view features for each depth candidate, thereby incurring high memory cost. To address the inherent memory issue, we introduce a Warp-Index Epipolar Attention mechanism that only stores warp indices for similarity matrix multiplication. We first compute the warping index map $\bm{I}$ as follows:
${\bm{I}}^{j\to i}=\mathcal{IW}({\bm{F}}^{j},{\bm{P}}^{i},{\bm{P}}^{j},\bm{G})\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times D},$ (4)
@@
Depth Probability Boosting Strategy. In each depth probability boosting unit, we stack $M$ Warp-Index Epipolar Attention layers to produce $M$ depth probability estimation results. To combine these isolated estimated outputs for stronger depth estimation capability, we propose a depth probability boosting strategy. Specifically, we initialize the depth probability matrix as $P_{0}=I$ and compute the subsequent updates as follows:
@@
Iterative Depth Estimation Process. Based on the depth probability boosting strategy, each Depth Unit produces an enhanced depth-probability map $P_{M}$. The residual depth map $\Delta\bm{D}$ is then computed by linearly combining $P_{M}$ with the corresponding depth-candidate vector $\bm{G}$:
${\Delta\bm{D}}={\bm{P}_{M}}{\bm{G}},$ (8)
where the combination is along the depth dimension. This residual depth $\Delta\bm{D}$ is added to the previous depth estimate to obtain a refined depth map at each iteration. We initialize the depth map as ${\bm{D}_{0}}=\bm{0}$, and the depth is then progressively computed in this manner:
@@
where $n\in\{1,\ldots,N\}$, and $\bm{D}_{N}$} is the final depth map of the iterative depth estimation process, from which the Gaussian mean parameters are obtained through unprojection. Notably, the depth search range is updated after each depth estimation unit. The search center is determined by the previous unit’s depth estimate, and the search range is halved. Thanks to our memory-efficient Warp-Index Epipolar Attention, IDESplat can progressively increase the feature resolution throughout this iterative process. Finally, IDESplat performs warping and similarity computation at the original input resolution of $256\times 256$, producing refined depth maps.

### v3

@@ Abstract
Generalizable 3D Gaussian Splatting aims to directly predict Gaussian parameters using a feed-forward network for scene reconstruction. Among these parameters, Gaussian means are particularly difficult to predict, so depth is usually estimated first and then unprojected to obtain the Gaussian sphere centers. Existing methods typically rely solely on a single warp to estimate depth probability, which hinders their ability to fully leverage cross-view geometric cues, resulting in unreliable and coarse depth maps. To address this limitation, we propose IDESplat, which iteratively applies warp operations to boost depth probability estimation for accurate Gaussian mean prediction. First, to eliminate the inherent unreliability of a single warp, we introduce a Depth Probability Boosting Unit (DPBU) that integrates epipolar attention maps produced by cascading warp operations in a multiplicative manner. Next, we construct an iterative depth estimation process by stacking multiple DPBUs, progressively identifying potential depth candidates with high likelihood. As IDESplat iteratively boosts depth probability estimates and updates the depth candidates, the depth map is gradually refined, resulting in accurate Gaussian means. We conduct experiments on RealEstate10K, ACID and DL3DV. IDESplat achieves outstanding reconstruction quality and state-of-the-art performance with real-time efficiency. On RE10K, it outperforms DepthSplat by 0.33 dB in PSNR, using only 10.7% of the parameters and 70% of the memory. Additionally, our IDESplat improves PSNR by 2.95 dB over DepthSplat on the DTU dataset in cross-dataset experiments, demonstrating its strong generalization ability. The code is available at https://github.com/CVL-UESTC/IDESplat.
@@ 1 Introduction
Efforts have been made to obtain accurate and refined depth estimation. Early methods [2, 34] directly use differentiable operations to predict the depth probability distribution from image features. Although these methods can predict the depth map for scene reconstruction in a generalizable way, their ability to exploit multi-view feature similarity is limited, which restricts their performance. Subsequent approaches [4, 20] introduce cost volumes that use warp operations to establish feature similarity across views, which provide valuable geometric cues for depth estimation and simplify the network learning process. However, these methods rely solely on a single warp to model feature similarity, which prevents them from fully exploiting the rich geometric information, leading to unreliable and coarse depth maps. Therefore, how to incorporate multiple warps to gradually leverage rich cross-view geometric details, producing refined and reliable depth maps for accurate Gaussian mean prediction, remains a key challenge.
In this paper, we propose IDESplat, which iteratively performs warps to refine depth maps for accurate Gaussian means prediction. By integrating cascade warp results, we can progressively boost the feature similarity measure to identify high-likelihood surface points and suppress low-probability depth candidates. The iterative warp framework of our IDESplat is shown in Fig. 1. Firstly, to eliminate the unreliability of a single warp, we introduce a Depth Probability Boosting Unit (DPBU) to fuse multiple epipolar attention maps in a multiplicative manner for a reliable depth map. Next, we gradually update the depth search range while increasing the feature resolution, enabling warp and correlation calculations at a finer scale. Feature matching becomes easier and more precise as the depth search range is re-centered and image features are enhanced in this process. Finally, for other Gaussian parameters, we propose a Gaussian Focused Module that determines the most relevant Gaussian tokens to compute attention weights for feature interaction. The experimental results on RealEstate10K show that IDESplat outperforms DepthSplat by 0.33 dB in PSNR with only 10.7% of the parameters and 70% of the memory. Moreover, IDESplat shows outstanding cross-dataset generalization, achieving 28.79 dB on ACID when transferred from RE10K, outperforming methods trained directly on ACID. It also improves PSNR by 2.95 dB over DepthSplat on the DTU dataset in cross-dataset experiments. In summary, the main contributions of this paper are as follows:
@@
  • To eliminate the inherent unreliability of a single warp, we introduce a Depth Probability Boosting Unit that multiplicatively integrates multiple epipolar attention maps for reliable and refined depth map estimation.
@@ 2 Related Work
Generalizable 3D Gaussian Splatting. Single-scene 3D Gaussian Splatting methods [16, 8, 48, 28, 9] enable more efficient rendering than neural fields and volume rendering methods [22, 31, 26, 41] thanks to their rasterization-friendly formulation, but they still suffer from long optimization time and limited generalization. To address this issue, generalizable 3D Gaussian Splatting methods [2, 35, 21, 46] predict all Gaussian parameters in a single feed-forward pass, enabling fast reconstruction and novel view synthesis for unseen scenes. Early works [2, 34] directly regressed Gaussian parameters from image features using feed-forward networks. With the introduction of cost-volume construction into generalizable 3DGS, subsequent methods [4, 20] exploited cross-view geometric cues to improve depth estimation and simplify Gaussian parameter prediction. More recently, methods [46, 21] further improved performance by incorporating pre-trained monocular depth models. However, existing methods still rely on a single warp for depth estimation, which limits their ability to fully exploit cross-view feature cues. In contrast, IDESplat integrates feature similarity from cascaded warps to produce more reliable and refined depth maps for Gaussian mean prediction.
Iteration-based Optimization Methods. Iteration-based optimization methods [1, 17] are widely used in various tasks [12, 56, 3] due to their ability to progressively enhance and refine the feature learning process. For instance, in optical flow estimation [36, 13, 49, 33, 14], iterative optimization is commonly used to build coarse-to-fine pyramidal features by stacking multiple feature units, resulting in more accurate and stable flow estimation. In the field of depth estimation, some works [18, 37, 38, 25, 42, 43, 6] have iteratively retrieved multi-view correlation features using structures like GRUs and LSTMs with shared parameters to update the disparity field, yielding better depth estimation results. Similarly, in monocular depth estimation [57, 30, 10], iterative methods are often used to gradually refine the depth search range, leading to more stable and accurate predictions. Additionally, in the 3D Gaussian Splatting (3DGS) domain, recent methods [39, 53, 44, 5, 11] have also attempted to design iterative optimization processes to reduce the difficulty of 3D Gaussian reconstruction tasks. Unlike existing iterative optimization methods, we propose a novel Depth Probability Estimation Unit that integrates the similarity results of multiple warps in a multiplicative manner. This approach produces more reliable and refined depth maps, while progressively enhancing feature resolution and refining the depth candidate range throughout the iterative process.
@@ 3 Method
Figure 2: The overall architecture of IDESplat. IDESplat consists of three key parts: a feature extraction backbone, an iterative depth probability estimation process, and a Gaussian Focused Module (GFM). The iterative process is built upon cascaded Depth Probability Boosting Units (DPBUs). Within each DPBU, depth probabilities are estimated sequentially by stacked Warp-Index Epipolar Attention (WIEA) blocks, where each block operates on progressively refined features and thus produces distinct probability estimates. Each unit combines multi-level warp results in a multiplicative manner to mitigate the inherent unreliability of a single warp. As IDESplat iteratively updates the depth candidates and boosts the probability estimates, the depth map becomes more precise, leading to accurate Gaussian means.
@@
Given a sequence of $V$ sparse-view images $\mathcal{I}=\{{\mathcal{I}}^{i}\}_{i=1}^{V}$, where each image ${\mathcal{I}}^{i}$ has dimensions $H\times W\times 3$, and the corresponding camera projection matrices are ${\mathcal{P}}^{i}\in\mathbb{R}^{3\times 4}$. The goal of the generalizable 3DGS task is to learn a network that maps images to 3D Gaussian parameters. This process is defined as:
$f_{\bm{\theta}}:\{({\mathcal{I}}^{i},{\mathcal{P}}^{i})\}_{i=1}^{V}\mapsto\{(\bm{\mu}_{j},\bm{\alpha}_{j},\bm{\Sigma}_{j},\bm{c}_{j})\}_{j=1}^{V\times H\times W},$ (1)
where $f_{\bm{\theta}}$ is a feed-forward network with learnable parameters $\bm{\theta}$}. The parameters $\bm{\mu}_{j}$, $\bm{\alpha}_{j}$, $\bm{\Sigma}_{j}$, and $\bm{c}_{j}$ are the Gaussian mean, opacity, covariance, and color, respectively. Since Gaussian means are difficult to predict directly, most existing methods estimate them by unprojecting a depth map into 3D. The depth estimation scheme is as follows: First, the input image sequence is processed using a multi-view feature extraction backbone, resulting in a downsampled feature $\bm{F}\in\mathbb{R}^{V\times\frac{H}{4}\times\frac{W}{4}\times C}$. Then, the cost volume is computed based on $\bm{F}$, the camera projection matrices ${\mathcal{P}}^{i}$, ${\mathcal{P}}^{j}$ for different views, and the depth candidates ${\bm{G}}=[d_{1},d_{2},\cdots,d_{D}]\in\mathbb{R}^{D}$. Specifically, the feature ${\bm{F}}^{j}$ from view $j$ is warped to view $i$ as follows:
${\bm{F}}^{j\to i}=\mathcal{W}({\bm{F}}^{j},{\mathcal{P}}^{i},{\mathcal{P}}^{j},\bm{G})\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times D\times C},$ (2)
@@
where $m\in\{1,\ldots,D\}$ and $C$ denotes the feature channel dimension. Subsequently, $\bm{C}^{i}=[\bm{C}_{d_{1}}^{i},\ldots,\bm{C}_{d_{D}}^{i}]$ is refined with a U-Net and upsampled to the input resolution as $\tilde{\bm{C}}^{i}\in\mathbb{R}^{H\times W\times D}$. Finally, a $\mathrm{softmax}$ function is applied to obtain the probability of each candidate depth, and the final depth map is computed as their weighted average.
To achieve high-quality 3D reconstruction, an accurate depth map $\bm{D}$ is crucial since it determines the centers of 3D Gaussians. However, this commonly used method only relies on a single cross-view warp to compute feature similarity, underutilizing cross-view geometry and often leading to unreliable, coarse depth estimates. In addition, this typical correlation computation method requires storing all the dense warping features, which incurs significant memory overhead, especially when the depth candidate size $D$ is large.
@@ 3.2 Iterative Multi-View Depth Estimation
Warp-Index Epipolar Attention. The warp operation is key to constructing the cost volume or epipolar attention map, as it establishes pixelwise correspondences between features in the source and target views. However, as shown in Eq. (2), this warping operation requires sampling target-view features for each depth candidate, thereby incurring high memory cost. To alleviate the inherent memory overhead, we introduce a Warp-Index Epipolar Attention mechanism that only stores warp indices for similarity matrix multiplication. We first compute the warping index map $\bm{I}$ as follows:
${\bm{I}}^{j\to i}=\mathcal{IW}({\bm{F}}^{j},{\mathcal{P}}^{i},{\mathcal{P}}^{j},\bm{G})\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times D},$ (4)
@@
Depth Probability Boosting Strategy. In each depth probability boosting unit, we stack $M$ Warp-Index Epipolar Attention layers to produce $M$ depth probability estimation results. To combine these isolated estimated outputs for stronger depth estimation capability, we propose a depth probability boosting strategy. Specifically, we initialize the depth probability matrix $\bm{P}_{0}$ as an all-ones matrix and compute the subsequent updates as follows:
@@
Iterative Depth Estimation Process. Based on the depth probability boosting strategy, each Depth Unit produces an enhanced depth-probability map $\bm{P}_{M,n}$ at iteration $n$. To refine depth estimates symmetrically around the current prediction, we formulate the depth update using a relative depth-candidate offset vector $\Delta\bm{G}_{n}$, which allows the network to predict both positive and negative residuals. Specifically, for the first iteration ($n=1$), we uniformly sample depth candidates within the initial range $d_{\min},d_{\max}]$. The depth-candidate vector used for feature matching is defined as $\bm{G}_{1}=[d_{\min}+{I}_{1},\ d_{\min}+2{I}_{1},\ \ldots,\ d_{\min}+\mathcal{D}_{1}{I}_{1}]$, where ${I}_{1}=(d_{\max}-d_{\min})/(\mathcal{D}_{1}+1)$ denotes the sampling interval and $\mathcal{D}_{1}$ is the number of candidates. Since the initial depth map is set to $\bm{D}_{0}=\bm{0}$, the relative offset vector in the first iteration is defined as $\Delta\bm{G}_{1}=\bm{G}_{1}$. For subsequent iterations ($n>1$), the depth search range is centered at the previous depth estimate $\bm{D}_{n-1}$. Accordingly, the depth-candidate vector is updated as $\bm{G}_{n}=[\bm{D}_{n-1}-k{I}_{n},\ \ldots,\ \bm{D}_{n-1},\ \ldots,\ \bm{D}_{n-1}+k{I}_{n}]$, and the corresponding relative offset vector is $\Delta\bm{G}_{n}=[-k{I}_{n},\ \ldots,\ 0,\ \ldots,\ k{I}_{n}]$. Here, the interval shrinks as ${I}_{n}={I}_{1}/n$, enabling progressively finer refinement over iterations, and the number of candidates is $\mathcal{D}_{n}=2k+1$. The residual depth map $\Delta\bm{D}_{n}$ is then computed as the weighted sum of the relative offsets $\Delta\bm{G}_{n}$ with the probability map $\bm{P}_{M,n}$ along the depth dimension:
${\Delta\bm{D}_{n}}={\bm{P}_{M,n}}{\Delta\bm{G}_{n}}.$ (8)
The refined depth map is updated additively at each iteration:
@@
where $n\in\{1,\ldots,N\}$. $\bm{D}_{N}$} is the final depth map of the iterative depth estimation process, from which the Gaussian mean parameters are obtained through unprojection. Enabled by our memory-efficient Warp-Index Epipolar Attention, IDESplat progressively increases the feature resolution over iterations. In the final stage, IDESplat performs warping and similarity computation at the original input resolution of $256\times 256$ to produce the refined depth map.